OpenStudy (anonymous):

Below are the data collected from two random samples of 500 American students on the number of hours they spend in school per day (rounded to the nearest hour): Meg concludes that students spend a mean of 7 hours in school each day. Tara thinks the mean is 6 hours. Who is correct—Meg or Tara? Help? I'm not exactly sure how to solve this.

OpenStudy (perl):

where is the data collected

OpenStudy (anonymous):

There it is (:

OpenStudy (perl):

to find the mean you have to multiply x by frequency

OpenStudy (perl):

sum (x * frequency of x ) / total sum

OpenStudy (anonymous):

Haha I'm kinda just sitting here trying to figure it out. Like, I found the mean of both data samples and they're both 100 but is that even useful information for this answer? Like, 5 hours is closest to the mean but that has nothing to do with it. . .

OpenStudy (perl):

not quite, the mean is 6

OpenStudy (perl):

sample A mean = ( 4 * 70 + 5*100 + 6*125 + 7*135 + 8 * 70 ) / ( 70 + 100 + 125 + 135 + 70)

OpenStudy (anonymous):

Ah, okay! I thought you were supposed to divide by the number of numbers? (sorry that hardly makes sense)

OpenStudy (perl):

If you are given just the raw ungrouped data, then you can add them up and divide by total. Here we are given grouped data in the form of a frequency distribution, so its a little different.

OpenStudy (perl):

so the first column for sample A is saying , there are 70 people who spend 4 hours a day in school. if you wanted to convert this to raw data you would have 4,4,4,4,4,4,4..... seventy fours

OpenStudy (perl):

so we have to modify the formula Sum x * f / sum x sample A= ( 4 * 70 + 5*100 + 6*125 + 7*135 + 8 * 70 ) / ( 70 + 100 + 125 + 135 + 70)

OpenStudy (anonymous):

ah okay! So I do the same for sample B?

OpenStudy (anonymous):

and then see which is the average?

OpenStudy (perl):

right